Documenta Mathematica

Extra Volume: John H. Coates’ Sixtieth Birthday, 2006

Preface  1
Foreword  3
Samegai’s Waters  5

K. Ardakov and K. A. Brown Ring-Theoretic Properties

of Iwasawa Algebras: A Survey  7–33

G. Banaszak, W. Gajda, P. Krasoń On the Image of l-Adic Galois Representations

for Abelian Varieties of Type I and II  35–75

Siegfried Böcherer^{∗}, A. A. Panchishkin^{†} Admissible p-adic Measures Attached to

Triple Products of Elliptic Cusp Forms  77–132

David Burns and Matthias Flach On the Equivariant Tamagawa Number Conjecture

for Tate Motives, Part II.  133–163

David Burns and Otmar Venjakob On the Leading Terms of Zeta Isomorphisms and

p-Adic L-functions in Non-Commutative Iwasawa Theory 165–209

Kevin Buzzard and Frank Calegari

The 2-adic Eigencurve is Proper.  211–232

L. Clozel, E. Ullmo Equidistribution Adélique des Tores

et Équidistribution des Points CM  233–260

Robert Coleman and Ken McMurdy

Fake CM and the Stable Model of X_{0}(Np^{3})  261–300

Daniel Delbourgo Λ-Adic Euler Characteristics

of Elliptic Curves  301–323

Ehud de Shalit Coleman Integration Versus

Schneider Integration on Semistable Curves  325–334

Ralph Greenberg On The Structure of

Certain Galois Cohomology Groups  335–391

Michael Harris, Jian-Shu Li, Christopher M. Skinner p-Adic L-Functions for Unitary Shimura Varieties

I: Construction of the Eisenstein Measure  393–464

Haruzo Hida

Anticyclotomic Main Conjectures  465–532

Frazer Jarvis Optimal Levels for Modular Mod 2

Representations over Totally Real Fields  533–550

Kazuya Kato Universal Norms of p-Units

in Some Non-Commutative Galois Extensions  551–565

Shinichi Kobayashi An Elementary Proof of the

Mazur-Tate-Teitelbaum Conjecture for Elliptic Curves 567–575

Barry Mazur, William Stein, John Tate

Computation of p-Adic Heights and Log Convergence  577–614

Robert Pollack and Tom Weston

Kida’s Formula and Congruences  615–630

P. Schneider, J. Teitelbaum Banach-Hecke Algebras

and p-Adic Galois Representations  631–684

Anthony J. Scholl

Higher Fields of Norms and (ϕ, Γ)-Modules  685–709

Joseph H. Silverman Divisibility Sequences and

Powers of Algebraic Integers  711–727

Richard Taylor On the Meromorphic Continuation

of Degree Two L-Functions  729–779

J. Tilouine

Siegel Varieties and p-Adic Siegel Modular Forms  781–817

J.-P. Wintenberger

On p-Adic Geometric Representations of G_{ℚ}  819–827